Waterway Company purchases equipment on January 1, Year 1, at a cost of $582,000. The asset is expected to have a service life of 12 years and a salvage value of $52,380.

1)Compute the amount of depreciation for each of Years 1 through 3 using the sum-of-the-years'-digits method.

2)Compute the amount of depreciation for each of Years 1 through 3 using the double-declining-balance method.
Answer #1

1) Sum of year digit = 78

Depreciable base = 582000-52380 = 529620

Depreciation expense
Year 1 529620*12/78 = 81480
Year 2 74690
Year 3 67900

2) Double decline rate = 100/12*2 = 16.67%

Depreciation expense
Year 1 582000/6 = 97000
Year 2 582000*5/6*1/6 = 80833
Year 3 582000*5/6*5/6*1/6 = 67361